one airplane holds 50 passengers more than another airplane. together they hold 450 passengers. how many passengers does each hold.
Answer by checkley77(12844)

x+(x+50)=450
2x+50=450
2x=450-50
2x=400
x=400/2
x=200 passengers for the smaller plane.
200+50=250 passengers for the larger plane.
Proof:
200+250=450
450=450